OpenStudy (anonymous):

Choose the ordered pair that satisfies 7x – 15y < -13.

OpenStudy (anonymous):

7x + 15y = 32 x = 3y + 20 you put x in for x. 7(3y+20)+15y=32 multiple by using the distrutive property. 21y+140+15y=32 add a like terms. 36y+140=32 take the opposite of 140(-140) and subtract it from both sides. 36y=-108 divide 36 to both sides. y=-3 then you fill the y in x=3y+20 x=3(-3)+20 multiple 3 times -3 x=-9+20 add -9 to 20 x=11

OpenStudy (anonymous):

here

OpenStudy (anonymous):

Thanks .

OpenStudy (anonymous):

im a nerd